9. ¿Cuál es el radio de la esfera dada por x2 + y2 + 22 - 10x +10y = 50?
Answer #1

00 x² + y2 +2² -lor + loy= So (x²-10x) + (y2 +164) +z? = 50 (x² - 10x +25-25) + lys +10y + 25-25) +2=50 (x-5) ² 2 + 3+5)* tz